OpenStudy (anonymous):

Find the circumference and the area of a circle with radius 5 yards. Use the value 3.14 as pie. Do not round your answers.Be sure to include the correct units in your answers.

OpenStudy (anonymous):

Do you know the formulas for the circumference and area of a circle?

OpenStudy (anonymous):

nope i just need the answers to both

OpenStudy (anonymous):

I'll give you the formulas, you can do the dirty work and plug the values in to solve it. \[C = 2\pi r\] \[A = \pi r ^{2}\]

OpenStudy (anonymous):

please can you just give me the answer this is the last prob i dont really want to do

OpenStudy (anonymous):

You need to learn it somehow.

OpenStudy (anonymous):

It's not hard.

OpenStudy (anonymous):

i know how to do but i always calculate it wrong please

OpenStudy (anonymous):

I would help you if I was convinced that you genuinely needed help. But I just gave you everything you need to know to solve it. Here, I'll define your variables as well so you can't get anything "mixed up": C = circumference A = area r = radius (half of the diameter) \[\pi = p i\] And pi = 3.14

OpenStudy (anonymous):

People shouldn't have to feed you the answers to your homework. You CAN do it on your own, you just have to put your mind to it.